Show that the Trigonometrical equationn Cos 42 + cos 78 + cos 162 = 0 ?

Grade:11

2 Answers

Mounik
8 years ago
The answer is take Lhs=
LHS = Cos 42 + cos 78 + cos 162
= cos(60 – 18)+ cos(60+18)+cos(180-18)
= cos 60 . cos 18 + sin 60 . sin 18 + cos 60 . cos 18 – sin60.sin 18 – cos 18
= 2cos 60 . cos 18 – cos 18
= 2. ½ . cos 18 – cos 18 
= 0 =RHS 
siva prasad reddy
8 years ago
LHS = Cos 42 + cos 78 + cos 162= cos(60 – 18)+ cos(60+18)+cos(180-18)= cos 60 . cos 18 + sin 60 . sin 18 + cos 60 . cos 18 – sin60.sin 18 – cos 18= 2cos 60 . cos 18 – cos 18= 2. ½ . cos 18 – cos 18 = 0 =RHS